OpenStudy (kainui):

What most adequately sums up the 90's?

OpenStudy (anonymous):

Is this a legit question?

OpenStudy (anonymous):

Really?

OpenStudy (kainui):

@ParthKohli says 90+91+92+93+...+99 but I'm not sure that's good enough.

OpenStudy (anonymous):

It is good enough baby.

OpenStudy (anonymous):

dude... you must be 1 20s boy

Parth (parthkohli):

\[\int_{90}^{100} xdx\]Memories, memories. ;-;

Parth (parthkohli):

If you define 90 to be in the 90s, then add 90 to that.

OpenStudy (shamil98):

\[\Large \sum_{i = 90}^{99} i\] same thing rite idk i suck at math

OpenStudy (kainui):

Yes but wait are these truly all the 90s? I mean not just the 90s that are integers are included which is good. But what about -92?

Parth (parthkohli):

Then the sum should be zero.

OpenStudy (shamil98):

if you included the negatives from -90 to -99 it would just be zero..

OpenStudy (kainui):

Yeah, so what is it? Also, I just realized something possibly important. \[\Large \pi ^2*10=98.696...\] Could this mean something we didn't consider?

OpenStudy (anonymous):

@ParthKohli I don't understand why I can get the same answer from your methods \[\int_{90}^{100} xdx =950\] while 90+91+....+99=945 why?

OpenStudy (texaschic101):

lol...I thought you were talking about the 90's like 1990 .....memories

OpenStudy (shamil98):

he's going from 90 to 100 we went from 90 to 99

Parth (parthkohli):

Because integrating is not summing up stepwise.

OpenStudy (shamil98):

or something like that

OpenStudy (kainui):

Well in fact there are an infinite number of 90s between 90 and 100, so we multiply each one by an infinitesimal amount and then add it up. So we're getting kind of like a "relative" sum or some such. Not sure. But the answer might be infinity or zero.

Parth (parthkohli):

Yeah, you would think that the 99 to 100 should be giving me a good advantage, but it doesn't. Know why?